Use x to represent my age in years
My brother is 5 years older than me
The sum of our ages is 111 years.

Write an equation using the information given and work out my age..

Answer by checkley79(3341)

x+(x+5)=111
2x=111-5
2x=106
x=106/2
x=53 age of the younger brother.
53+5=58 age of the older brother.
Proof:
53+58=111
111=111